OpenStudy (eric_d):

Expand (1+x)^-1/4 in ascending powers of x up to the term in x^2. Prove that 3/2(1+1/80)^-1/4=5^1/4. Using your expansion for (1+x)^-1/4 and x=1/80,find an approximation for 5^1/4, giving yr answer correct to 5d.p.

OpenStudy (eric_d):

@ganeshie8

OpenStudy (eric_d):

@ikram002p

OpenStudy (ikram002p):

hehe i will try what ganesh did in the previouse qn else i have no idea :P

OpenStudy (eric_d):

ok

ganeshie8 (ganeshie8):

yeah binomial expansion again

ganeshie8 (ganeshie8):

start by writing out the binomial expansion formula

ganeshie8 (ganeshie8):

\[\large \begin{align} (x+1)^{-1/4} &= \sum \limits_{k=0}^{\infty } \binom{-1/4}{k} x^r \\ &=\binom{-1/4}{0} x^0 +\binom{-1/4}{1} x^1 +\binom{-1/4}{2} x^2 + \cdots \end{align} \]

ganeshie8 (ganeshie8):

@ikram002p wana practice binomial expansion with fractions ? :)

OpenStudy (ikram002p):

yeah , but i have to go nw >.< will you teach me this ?

ganeshie8 (ganeshie8):

really nothing much to teach you lol you already know binomial expansion, so simply extend it to fractions and hope that it works, thats all :)

OpenStudy (ikram002p):

oki , ill try when i come back mmmm

OpenStudy (eric_d):

Thanks @ganeshie8

ganeshie8 (ganeshie8):

np :) have u finished it ?